# New Skyrme parametrizations to describe finite nuclei and neutron star matter with realistic effective masses. II. Adjusting the spin-dependent terms

### v1

@@ Abstract
Many common Skyrme functionals present ferromagnetic instabilities or unrealistic density dependence of the spin-1 Landau parameters. To solve these problems, we consider the Skyrme interaction as a density-functional rather than a density-dependent two-body force. This allows us to adjust the spin-dependent terms of the new extended Skyrme functionals of our previous paper [M. Duan and M. Urban, Phys. Rev. C 110, 065806 (2024)] independently without altering the properties of spin saturated matter. The parameters of the spin-dependent terms are determined by fitting the Landau parameters $G_{0}$ and $G_{0}^{\prime}$ in neutron matter and symmetric nuclear matter and the effective-mass splitting of up and down particles in spin polarized matter to the results of microscopic calculations. Using the new parametrizations, called Sky3s and Sky4s, the spin-related properties of nuclear matter are in good agreement with the microscopic results. As an application, we compute response functions and neutrino scattering rates of neutron-star matter with the new functionals having realistic effective masses and Landau parameters.
@@ I Introduction
An important application of the nuclear energy density functional is to describe the transport properties of neutrinos in (proto-)neutron star and supernova matter. Neutrinos play a crucial role in explaining supernova explosions, neutron star mergers, proto-neutron star evolution, and neutron star cooling (Hirata et al., 1987; Mezzacappa, 2005; Janka, 2012; Burrows, 2013; Burrows and Lattimer, 1986; Mayle et al., 1987; Keil, 1994; Pascal et al., 2022; Yakovlev and Pethick, 2004). These astrophysical processes occur along with neutrino absorption and scattering. As many studies have done, such as the early literature Iwamoto and Pethick (1982) and the recent references Pastore et al. (2015); Oertel et al. (2020); Duan and Urban (2023), the neutrino rates (absorption and scattering rates) or the neutrino mean free path can be related to the response functions, including responses computed using the Skyrme interactions.
@@
A couple of articles have studied neutrino mean free path using Skyrme energy density functional for pure neutron matter (PNM) (Pastore et al., 2012, 2014, 2015) and asymmetric nuclear matter (Davesne et al., 2014, 2019). But we recently reported that many Skyrme effective interactions predict an unrealistic density dependence of the effective masses, resulting, e.g., in the unphysical feature that the neutron Fermi velocity exceeds the speed of light at relatively low densities (Duan and Urban, 2023). Therefore new parametrizations were needed to solve this problem. We constructed two new Skyrme interactions, Sky3 and Sky4, solving the problem by adjusting the parameters to Brueckner-Hartree-Fock (BHF) results for the effective masses (Duan and Urban, 2024).
@@
However, there is still a problem in these two interactions, namely, they predict a ferromagnetic instability at densities that may be realized in neutron stars (Duan and Urban, 2024). The reason is that the spin dependent terms of the functionals are expressed in terms of the same parameters as the spin-independent ones, and the fitting focuses on observables that depend only on the spin-independent terms with the exception of the spin-orbit parameter $W_{0}$. Apart from the unphysical instability, the spin dependent terms affect also the neutrino scattering rates which involve contributions of spin-0 (spin-independent) and spin-1 (spin-dependent) channels. The response functions of neutron-star and supernova matter, thus the neutrino rates, can be computed only if the Landau parameters in the spin-1 channel have the correct values. Therefore, the spin-dependent terms of the new Skyrme functionals should be constrained reasonably. This is possible if one interprets the Skyrme interaction as a density functional, in which the spin-independent and the spin-dependent terms can be adjusted independently Bender et al. (2002).
@@
In this work, we first point out the problems of the spin-related properties of nuclear matter with the unconstrained Skyrme functionals Sky3 and Sky4, but also with other Skyrme functionals, in Sec. II. Then, we present the strategies to adjust the spin-dependent terms of the functionals by using results of BHF calculations in Sec. III. The Landau parameters in pure neutron matter and symmetric nuclear matter (SNM), the equation of state of spin-polarized neutron matter, and the instabilities in neutron star matter computed with the resulting new functionals Sky3s and Sky4s are also shown in Sec. III. Finally, we present some results of the full RPA (random phase approximation) response functions of neutron star matter and the corresponding neutrino scattering rates computed with Sky3s in Sec. IV, and summarize in Sec. V.
@@ II Spin-related properties of nuclear matter with Skyrme functionals
The standard form of the Skyrme energy density functional consists of the kinetic-energy term, the zero-range term, the density-dependent term, the effective-mass term, the finite-range term, the spin-orbit term, the tensor coupling term, and the Coulomb energy density (Chabanat et al., 1997). The actual expressions can vary from one parametrization to another. For example, density-dependent generalizations of the non-local ($t_{1}$ and $t_{2}$) terms of Skyrme interactions have been added into the BSk functionals (Chamel et al., 2009). In interactions developed for astrophysical applications, the tensor coupling term is usually omitted (Goriely et al., 2010, 2013, 2016). The KIDS interactions have generalized the density dependent ($t_{3}$) term to several $t_{3i}$ terms with corresponding exponents $\alpha_{i}$ (Gil et al., 2019). In our previous work (Duan and Urban, 2024), the functionals Sky3 and Sky4 have been constructed combining the modification of the density-dependent term with the density-dependent generalizations of the $t_{1}$ and $t_{2}$ terms.
According to Refs. Bender et al. (2003); Duan and Urban (2023), the spin-dependent parts of the Skyrme energy density functional can be written using standard notations as
@@
The parameter $\eta_{J}$, in the notation of Bender et al. (2003), indicates whether the contribution of the $t_{1}$ and $t_{2}$ terms to the spin-orbit field should be included ($\eta_{J}=1$) or not ($\eta_{J}=0$). For the relation between the spin current $\mathbbm{J}$ and the vector spin current $\bm{\mathrm{J}}$ see e.g. Ref. Lesinski et al. (2007).
Figure 1: The Landau parameters $G_{0}$ in pure neutron matter (top panel) and $G_{0}$ and $G_{0}^{\prime}$ in symmetric nuclear matter (middle and bottom panels). The results are computed with Skyrme interactions Sky3, Sky4 (Duan and Urban, 2024), SLy4 (Chabanat et al., 1998), BSk20 (Goriely et al., 2010), SLy5st (Margueron and Sagawa, 2009) and BSk17st (Margueron et al., 2009). The square symbols represent the microscopic BHF results of Refs. Zuo et al. (2003a, b).
@@
The Landau parameters computed with Sky3 and Sky4 were already shown in Ref. Duan and Urban (2024). We display them again in Fig. 1 together with the results computed with other Skyrme interactions, such as SLy4 (Chabanat et al., 1998), BSk20 (Goriely et al., 2010), SLy5st (Margueron and Sagawa, 2009), and BSk17st (Margueron et al., 2009). SLy4 and BSk20 are, respectively, examples for standard and extended Skyrme interactions. SLy5st and BSk17st represent Skyrme interactions whose spin-dependent parts have been modified in an attempt to constrain them using BHF results (Margueron and Sagawa, 2009; Margueron et al., 2009). For comparison, the microscopic BHF results (Zuo et al., 2003a, b)  [^1] From Fig. 2 of Ref. Zuo et al. (2003) one can infer that there is a typo in Table 1 of Ref. Zuo et al. (2003a) and $G_{0}$ in pure neutron matter for $\rho=0.8\;\text{fm}^{-3}$ should read $3.52$ instead of $2.52$, but probably this value is still incorrect, cf. discussion below Eq. (13). are shown as square symbols. As we can see, the results computed with Sky3 and Sky4 are nearest to the BHF values at saturation density, but their evolutions with density are very different from the BHF results. Other Skyrme interactions present not only different evolutions but also different values at saturation density, compared with BHF results. We conclude that the fitting of the parameters of the considered interactions is not sufficient to make them predictive in the spin-1 channel. And even with the additional parameters introduced in SLy5st (Margueron and Sagawa, 2009) and BSk17st (Margueron et al., 2009) there is not enough freedom to reproduce the BHF results.
@@
In the case of Sky3 and Sky4, the situation is worse because above some critical densities they give $G_{0}<-1$, violating the Landau stability criterion. To further illustrate this instability problem, we show in Fig. 2 the equations of state (EoS) of fully spin-polarized and unpolarized neutron matter computed with Sky3 and Sky4. For both Sky3 and Sky4, the fully polarized neutron matter becomes energetically favorable over the unpolarized one at densities above the respective critical densities indicated by the red and steel-blue dotted lines. The fundamental cause of this behaviour is the wrong values of the Landau parameters mentioned above. Otherwise, one would get a stiffer EoS for fully polarized than for unpolarized neutron matter. Therefore, the solution is to revise Sky3 and Sky4 to obtain realistic values of the Landau parameters.
@@ III.1 Determination of the spin-dependent terms of the new Skyrme functionals
We now focus on constraining the remaining coefficients $C_{0}^{s}$, $C_{1}^{s}$, $C_{0}^{sT}$, and $C_{1}^{sT}$. Notice that the choice $\eta_{J}=0$ implies that, in Sky3 and Sky4, we have $C_{0}^{sT}=C_{1}^{sT}=0$. This means that in spin-polarized matter, the spin up and down nucleons have the same effective mass, since
@@
Then $C^{sT}_{00}$ and $C^{sT}_{01}$ can be obtained by fitting the splitting between the up and down nucleon effective masses for spin-polarized symmetric nuclear matter, while $C^{sT}_{00}+C^{sT}_{10}$ and $C^{sT}_{01}+C^{sT}_{11}$ can be obtained by fitting the splitting between the up and down neutron effective masses for spin-polarized neutron matter. To get information about the density dependence, the results published in Refs. Bombaci et al. (2006); Bigdeli et al. (2009) are not enough and we obtained more data at different densities directly from the authors of these papers Bigdeli and Bordbar (2024); Vidaña (2024). We show $C^{sT}$ as functions of density in pure neutron matter and symmetric nuclear matter in Fig. 3. As we can see, $C^{sT}_{I}$ cannot be considered constant.
Besides, as mentioned above, the $C_{I}^{s}$ and $C_{I}^{sT}$ coefficients can be related to the Landau parameters $G_{0}$ and $G_{0}^{\prime}$ in pure neutron matter and symmetric nuclear matter. Therefore, we are now concentrating on obtaining realistic values of the Landau parameters for our two functionals. As we did above for the $C^{sT}$ coefficients and in Ref. Duan and Urban (2024) for the effective masses, we will again use the results of microscopic BHF calculations for this purpose.
@@
Hence, only $C^{s}_{00}$ and $C^{s}_{10}$ are defined by the values of $t_{0}$ and $y_{0}$ of Sky3 or Sky4, respectively. Because we suspect that there was a mistake in the calculation of $G_{0}$ in pure neutron matter in Refs.Zuo et al. (2003a, 2003) (see below), we determine $C^{s}_{I1}\dots C^{s}_{I3}$ by fitting only the Landau parameters $G_{0}$ and $G_{0}^{\prime}$ in symmetric nuclear matter to those computed using the microscopic BHF theory in Ref. Zuo et al. (2003b). According to Zuo et al. (2003), $G_{0}$ in pure neutron matter is computed from the spin susceptibility $\chi$ as
@@
where $\chi_{F}$ is the spin susceptibility of an ideal gas. By analyzing the different results for $G_{0}$ and $\chi$ shown in Ref. Zuo et al. (2003), one finds that in the calculation with three-body force, the factor $\frac{m^{*}}{m}$ in Eq. (13) was by mistake taken to be the same as in the calculation without three-body force. We can compute corrected values of $G_{0}$ from Eq. (13) using the results for $\tfrac{\chi}{\chi_{F}}$ given in Zuo et al. (2003) and for $\tfrac{m^{*}}{m}$ from Baldo et al. (2014) (which are similar to the Sky3 and Sky4 ones). We use only the last point (at $\rho=0.8\;\text{fm}^{-3}$) of the corrected $G_{0}$ to roughly guide the high-density behavior of the fits. The results of these fits are listed in Table 1.

### v2

@@ Abstract
Many common Skyrme functionals present ferromagnetic instabilities or unrealistic density dependence of the spin-dependent Landau parameters. To solve these problems, we consider the Skyrme interaction as a density-functional rather than a density-dependent two-body force. This allows us to adjust the spin-dependent terms of the new extended Skyrme functionals of our previous paper [M. Duan and M. Urban, Phys. Rev. C 110, 065806 (2024)] independently without altering the properties of spin saturated matter. The parameters of the spin-dependent terms are determined by fitting the Landau parameters $G_{0}$ and $G_{0}^{\prime}$ in neutron matter and symmetric nuclear matter and the effective-mass splitting of up and down particles in spin polarized matter to the results of microscopic calculations. Using the new parametrizations, called Sky3s and Sky4s, the spin-related properties of nuclear matter are in good agreement with the microscopic results. As an application, we compute response functions and neutrino scattering rates of neutron-star matter with the new functionals having realistic effective masses and Landau parameters.
@@ I Introduction
An important application of the nuclear energy density functional is to describe the transport properties of neutrinos in (proto-)neutron star and supernova matter. Neutrinos play a crucial role in explaining supernova explosions, neutron star mergers, proto-neutron star evolution, and neutron star cooling (Hirata et al., 1987; Mezzacappa, 2005; Janka, 2012; Burrows, 2013; Burrows and Lattimer, 1986; Mayle et al., 1987; Keil, 1994; Pascal et al., 2022; Yakovlev and Pethick, 2004). These astrophysical processes occur along with neutrino absorption and scattering. As many studies have done, such as the early literature Iwamoto and Pethick (1982) and the recent references Pastore et al. (2015); Oertel et al. (2020); Duan and Urban (2023); Lin et al. (2023), the neutrino rates (absorption and scattering rates) or the neutrino mean free path can be related to the response functions, including responses computed using the Skyrme interactions.
@@
A couple of articles have studied neutrino mean free path using Skyrme energy density functional for pure neutron matter (Pastore et al., 2012, 2014, 2015) and asymmetric nuclear matter (Davesne et al., 2014, 2019). But we recently reported that many Skyrme effective interactions predict an unrealistic density dependence of the effective masses, resulting, e.g., in the unphysical feature that the neutron Fermi velocity exceeds the speed of light at relatively low densities (Duan and Urban, 2023). Therefore new parametrizations were needed to solve this problem. We constructed two new Skyrme interactions, Sky3 and Sky4, solving the problem by adjusting the parameters to Brueckner-Hartree-Fock (BHF) results for the effective masses (Duan and Urban, 2024). Figure 1 displays the effective masses in pure neutron matter (PNM) and in symmetric nuclear matter (SNM) from the BHF calculations of Ref. Baldo et al. (2014) and the corresponding ones of Sky3 (which are identical to those of Sky4). We can see that the neutron effective mass in PNM and the nucleon effective mass in SNM computed with Sky3 and Sky4, which will be used in the calculations and fits in the present paper, are in good agreement with the BHF ones, although some discrepancies cannot be avoided, mainly because the relation $1/m_{p}^{*}(\text{PNM})+1/m_{n}^{*}(\text{PNM})=2/m^{*}(\text{SNM})$ imposed by Skyrme functionals is not satisfied by the BHF effective masses Duan and Urban (2024).
Figure 1: (Top) Neutron effective mass in PNM, (middle) proton effective mass in PNM (= isovector effective mass), and (bottom) nucleon effective mass in SNM (= isoscalar effective mass) from the BHF calculations of Ref. Baldo et al. (2014) (dash-dot lines) and the corresponding ones of Sky3 and Sky4 Duan and Urban (2024) (solid lines).
@@
However, there is still a problem in these two interactions, namely, they predict a ferromagnetic instability at densities that may be realized in neutron stars (Duan and Urban, 2024). The reason is that the spin dependent terms of the functionals are expressed in terms of the same parameters as the spin-independent ones, and the fitting focuses on observables that depend only on the spin-independent terms with the exception of the spin-orbit parameter $W_{0}$. Apart from the unphysical instability, the spin dependent terms affect also the neutrino scattering rates which involve contributions of spin-0 (spin-independent) and spin-1 (spin-dependent) channels. The response functions of neutron-star and supernova matter, thus the neutrino rates, can be computed only if the Landau parameters in the spin-1 channel have the correct values. Therefore, the spin-dependent terms of the new Skyrme functionals should be constrained reasonably.
The spin Landau parameters can be related to spin- and spin-isospin excitations, such as magnetic dipole (M1) and Gamow-Teller (GT) resonances. Already more than 40 years ago, the spin dependent Landau parameters at saturation density have been obtained from fits to GT-resonance systematics Bertsch et al. (1981); Gaarde et al. (1981); Suzuki (1982); Bertsch (1981). Research on this topic has been pursued since then Bender et al. (2002); Wakasa et al. (2005, 2012); Yasuda et al. (2018). Thus, a possibility to constrain the spin properties of Skyrme functionals (especially in the isovector channel) was to include data for Gamow-Teller resonances Osterfeld (1992); Ichimura et al. (2006) into the fitting protocol. Examples for interactions constrained in this way are SGII Giai and Sagawa (1981) and SAMi Roca-Maza et al. (2012). However, in these interactions the spin dependent and the spin independent terms are linked to each other, since they are all determined by the standard Skyrme parameters. A different approach was taken in Ref. Bender et al. (2002). There, the Skyrme interaction was interpreted as a density functional instead of a density-dependent two-body force, which allowed the authors to adjust the spin dependent terms independently to describe the Gamow-Teller strength distribution, without changing the other terms (which in that work were taken from the SkO parametrization Reinhard et al. (1999)). In the present work, we will also adjust the spin dependent terms independently, but instead of fitting them to Gamow-Teller resonances, we will adjust them to the results of microscopic calculations Zuo et al. (2003, 2003a, 2003b); Bigdeli et al. (2009); Vidaña et al. (2016); Bigdeli and Bordbar (2024); Vidaña (2024), similar to earlier attempts to improve the spin properties of the SLy5 and BSk17 interactions Margueron and Sagawa (2009); Margueron et al. (2009).
@@
In Sec. II, we first point out the problems of the spin-related properties of nuclear matter with the unconstrained Skyrme functionals Sky3 and Sky4, but also with other Skyrme functionals. Then, we present the strategies to adjust the spin-dependent terms of the functionals by using results of BHF calculations in Sec. III. The Landau parameters in pure neutron matter and symmetric nuclear matter, the equation of state of spin-polarized neutron matter, and the instabilities in neutron star matter computed with the resulting new functionals Sky3s and Sky4s are also shown in Sec. III. Finally, we present some results of the full RPA (random phase approximation) response functions of neutron star matter and the corresponding neutrino scattering rates computed with Sky3s in Sec. IV, and summarize in Sec. V.
@@ II Spin-related properties of nuclear matter with Skyrme functionals
The standard form of the Skyrme energy density functional consists of the kinetic-energy term, the zero-range term, the density-dependent term, the effective-mass term, the finite-range term, the spin-orbit term, the tensor coupling term, and the Coulomb energy density (Chabanat et al., 1997). The actual expressions can vary from one parametrization to another. For example, density-dependent generalizations of the non-local ($t_{1}$ and $t_{2}$) terms of Skyrme interactions have been added into the BSk functionals (Chamel et al., 2009). In interactions developed for astrophysical applications, the tensor coupling term is usually omitted (Goriely et al., 2010, 2013, 2016). The KIDS interactions have generalized the density dependent ($t_{3}$) term to several $t_{3i}$ terms with corresponding exponents $\alpha_{i}$ (Gil et al., 2019). In our previous work (Duan and Urban, 2024), the functionals Sky3 and Sky4 have been constructed combining the modification of the density-dependent term with the density-dependent generalizations of the $t_{1}$ and $t_{2}$ terms, namely, the functionals contain not only the $t_{3i}$ terms but also the density-dependent non-local $t_{4}$ and $t_{5}$ terms as in the extended BSk functionals.
As mentioned in Refs. Chamel et al. (2009); Duan and Urban (2024), the $(\bm{\mathrm{\nabla}}\rho)^{2}$ terms cannot be transformed into $\rho\Delta\rho$ by integration by parts in the extended Skyrme functionals, in contrast to standard Skyrme functionals. Similarly, also the $(\bm{\mathrm{\nabla}}\otimes\bm{\mathrm{s}})^{2}=\sum_{ij}(\nabla_{i}s_{j})^{2}$ terms cannot be transformed into $\bm{\mathrm{s}}\cdot\Delta\bm{\mathrm{s}}$ by integration by parts Pastore et al. (2015); Duan and Urban (2024). Therefore, within the form chosen in Ref. Duan and Urban (2024), the spin-dependent parts of the Skyrme energy density functional can be written using standard notations as Bender et al. (2003); Duan and Urban (2023)
@@
The parameter $\eta_{J}$, in the notation of Bender et al. (2003), indicates whether the contribution of the $t_{1}$ and $t_{2}$ terms to the spin-orbit field should be included ($\eta_{J}=1$) or not ($\eta_{J}=0$). In our new functionals Sky3 and Sky4, we use $\eta_{J}=0$, i.e., we have $C^{sT}_{0}=C^{sT}_{1}=0$ in these two functionals. For the relation between the spin current $\mathbbm{J}$ and the vector spin current $\bm{\mathrm{J}}$ see e.g. Ref. Lesinski et al. (2007).
Concerning the density dependence of the $C$ coefficients, in Ref. Duan and Urban (2024), the choice was made to write them as polynomials of $\rho^{1/3}$ (corresponding to an expansion in powers of $k_{F}$). Hence, in Sky3 and Sky4, the exponents $\alpha_{i}$, $\beta$, and $\gamma$ were not fitted. In the $t_{3}$ and $y_{3}$ terms, they are $\alpha_{i}=i/3$, as in Ref. Gil et al. (2019). Limiting the density dependence of the $C^{\rho}$ coefficients (i.e., the $t_{3}$ and $y_{3}$ terms) to order $\rho$ ($i=3$) allowed us to precisely fit microscopically computed PNM EoS from extremely low to the highest densities. In order not to interfere with the EoS fit, the density dependence of the effective mass term was limited to order $\rho^{1/3}$ as explained in Ref. Duan and Urban (2024), and hence the exponents are $\beta=\gamma=1/3$ in Sky3 and Sky4, while they have different values in the extended Skyrme forces of the BSk family Chamel et al. (2009).
Figure 2: The Landau parameters $G_{0}$ in pure neutron matter (top panel) and $G_{0}$ and $G_{0}^{\prime}$ in symmetric nuclear matter (middle and bottom panels). The results are computed with Skyrme interactions Sky3, Sky4 (Duan and Urban, 2024), SLy4 (Chabanat et al., 1998), BSk20 (Goriely et al., 2010), SLy5st (Margueron and Sagawa, 2009), BSk17st (Margueron et al., 2009), and SAMi Roca-Maza et al. (2012). The square symbols represent the microscopic BHF results of Refs. Zuo et al. (2003a, 2003, b).
@@
The Landau parameters computed with Sky3 and Sky4 were already shown in Ref. Duan and Urban (2024). We display them again in Fig. 2 together with the results computed with other Skyrme interactions, such as SLy4 (Chabanat et al., 1998), BSk20 (Goriely et al., 2010), SLy5st (Margueron and Sagawa, 2009), BSk17st (Margueron et al., 2009), and SAMi Roca-Maza et al. (2012). SLy4 and BSk20 are, respectively, examples for standard and extended Skyrme interactions. SLy5st and BSk17st represent Skyrme interactions whose spin-dependent parts have been modified in an attempt to constrain them using BHF results (Margueron and Sagawa, 2009; Margueron et al., 2009). SAMi is a standard Skyrme force but fitted with special attention to Gamow-Teller resonances. For comparison, the microscopic BHF results (Zuo et al., 2003a, b)  [^1] There is a typo in Table 1 of Ref. Zuo et al. (2003a) and $G_{0}$ in pure neutron matter for $\rho=0.8\;\text{fm}^{-3}$ should read $3.52$ instead of $2.52$ Zuo (2025). are shown as square symbols. As we can see, the results computed with Sky3 and Sky4 are close to the BHF values at saturation density, but their evolutions with density are very different from the BHF results. Other Skyrme interactions present not only different evolutions but also different values at saturation density. Of course, it is not a necessary condition that a reasonable functional must reproduce the BHF results, that in addition do not have a quantified uncertainty. But it is interesting that even in the cases of SLy5st (Margueron and Sagawa, 2009) and BSk17st (Margueron et al., 2009), which used the BHF results for the adjustment of their additional parameters, there is not enough freedom to reproduce the BHF results.
@@
In the case of Sky3, Sky4, and SAMi, the situation is worse because above some critical densities they give $G_{0}<-1$, violating the Landau stability criterion. To further illustrate this instability problem, we show in Fig. 3 the equations of state (EoS) of fully spin-polarized and unpolarized neutron matter computed with Sky3 and Sky4. For both Sky3 and Sky4, the fully polarized neutron matter becomes energetically favorable over the unpolarized one at densities above the respective critical densities indicated by the red and steel-blue dotted lines. The fundamental cause of this behaviour is the wrong values of the Landau parameters mentioned above. Otherwise, one would get a stiffer EoS for fully polarized than for unpolarized neutron matter. Therefore, the solution is to revise Sky3 and Sky4 to obtain realistic values of the Landau parameters.
@@ III.1 Determination of the spin-dependent terms of the new Skyrme functionals
We now focus on constraining the remaining coefficients $C_{0}^{s}$, $C_{1}^{s}$, $C_{0}^{sT}$, and $C_{1}^{sT}$. As mentioned before, the choice $\eta_{J}=0$ implies that, in Sky3 and Sky4, we have $C_{0}^{sT}=C_{1}^{sT}=0$. This means that in spin-polarized matter, the spin up and down nucleons have the same effective mass, since
@@
Then $C^{sT}_{00}$ and $C^{sT}_{01}$ can be obtained by fitting the splitting between the up and down nucleon effective masses for spin-polarized symmetric nuclear matter, while $C^{sT}_{00}+C^{sT}_{10}$ and $C^{sT}_{01}+C^{sT}_{11}$ can be obtained by fitting the splitting between the up and down neutron effective masses for spin-polarized neutron matter. To get information about the density dependence, the results published in Refs. Bombaci et al. (2006); Bigdeli et al. (2009) are not enough and we obtained more data at different densities directly from the authors of these papers Bigdeli and Bordbar (2024); Vidaña (2024).
The results computed by Ref. Bigdeli and Bordbar (2024) used only the Argonne V18 two-body force, while Ref. Vidaña (2024) also added the effect of the phenomenological three-body force (TBF) Urbana IX (reduced to a density dependent two-body force). Using these results is slightly inconsistent because the Sky3 and Sky4 effective masses were adjusted to the results of Ref. Baldo et al. (2014) obtained with a microscopically derived TBF Grangé et al. (1989), which was also employed in the calculation of the Landau parameters Zuo et al. (2003a, 2003, b). While the TBF has almost no effect at low density, it causes large uncertainties at high density. However, we have no choice because to our knowledge there are no results for the splitting of up and down effective masses computed with AV18 and the microscopic TBF in the present literature, and recomputing it using a microscopic TBF is a difficult task that clearly goes beyond the scope of this work. Moreover, our purpose is to get realistic spin properties instead of reproducing all nuclear matter properties computed using the microscopic TBF. Therefore, this imperfection should be acceptable for now.
In Fig. 4, we show $C^{sT}$ as functions of density in pure neutron matter and symmetric nuclear matter. As we can see, $C^{sT}_{I}$ cannot be considered constant. Obviously, the given fit is not the only possible one, but since we are fitting our parameters to theoretical results whose uncertainties are unknown, it is difficult to give a quantitative estimate of the uncertainties of the fitted $C^{sT}$ coefficients.
As mentioned above, the $C_{I}^{s}$ and $C_{I}^{sT}$ coefficients can be related to the Landau parameters $G_{0}$ and $G_{0}^{\prime}$ in pure neutron matter and symmetric nuclear matter. Therefore, the next step is to obtain realistic values of the Landau parameters for our two functionals. As we did above for the $C^{sT}$ coefficients and in Ref. Duan and Urban (2024) for the effective masses, we will again use the results of microscopic BHF calculations for this purpose.
@@
Hence, only $C^{s}_{00}$ and $C^{s}_{10}$ are defined by the values of $t_{0}$ and $y_{0}$ of Sky3 or Sky4, respectively. We determine $C^{s}_{I1}\dots C^{s}_{I3}$ by fitting only the Landau parameters $G_{0}$ and $G_{0}^{\prime}$ in symmetric nuclear matter to those computed using the microscopic BHF theory in Ref. Zuo et al. (2003b), because we found a problem while analyzing the results for pure neutron matter in Refs. Zuo et al. (2003a, 2003) (the results presented in these two references are the same), as described below. According to Ref. Zuo et al. (2003), $G_{0}$ in pure neutron matter is computed from the spin susceptibility $\chi$ as
@@
where $\chi_{F}$ is the spin susceptibility of an ideal gas. By analyzing the different results for $G_{0}$ and $\chi$ shown in Refs. Zuo et al. (2003a, 2003), one finds that in the calculation with TBF, the values of the factor $\frac{m^{*}}{m}$ are almost the same as those computed without TBF. It is worth noting that Refs. Zuo et al. (2003a, 2003) give almost the same $\frac{m^{*}}{m}$ as Ref. Baldo et al. (2014) in the two-body force case, while they give very different results when including the TBF, since the TBF seems to have almost no effect on $\frac{m^{*}}{m}$ in Refs. Zuo et al. (2003a, 2003) while it has a strong effect in Ref. Baldo et al. (2014). Since the Sky3 and Sky4 effective masses in PNM are similar to the results from Ref. Baldo et al. (2014) (to which they were adjusted) instead of those from Refs. Zuo et al. (2003a, 2003), we should not fit $G_{0}$ of Refs. Zuo et al. (2003a, 2003) directly, because it is normalized with $N_{0}=\frac{m^{*}k_{F}}{\pi^{2}\hbar^{2}}$ that contains the neutron effective mass. However, we can fit $\tfrac{\chi}{\chi_{F}}$ in essence. Therefore, we recompute $G_{0}$ from Eq. (13) using the results for $\tfrac{\chi}{\chi_{F}}$ given in Refs. Zuo et al. (2003a, 2003) and for $\tfrac{m^{*}}{m}$ from Ref. Baldo et al. (2014). Let us mention that in the case of SNM, this problem does not exist, because the ratios $\frac{m^{*}}{m}$ that can be deduced from the SNM results given in Ref. Zuo et al. (2003b) agree well with those of Sky3 and Sky4. The results of the fits are listed in Table 1.
Again, as in the case of the $C^{sT}$ coefficients, since we are fitting to theoretical calculations with unknown uncertainties, we cannot provide errors estimates for the fitted parameters. Just from comparing the results for the Landau parameters obtained in Refs. Zuo et al. (2003a, 2003, b) with and without TBF and considering the large uncertainties of the TBF, it is clear that there may be substantial errors in the Landau parameters.